OpenStudy (anonymous):

Which of the following is equal to sqrt(50a^6b^7)? a)5a^3b^3 sqrt(2b) b) 5a^3b^2 sqrt(2b) c) 10a^3b^3 d) 5^3b^3

OpenStudy (tkhunny):

Are you sitting in your exam right now?

OpenStudy (anonymous):

unforunately. @tkhunny

OpenStudy (tkhunny):

Why should anyone help you with that? Do we KNOW that a > 0? That's important.

OpenStudy (anonymous):

just because i need help during this exam and i posted here doesn't mean i'm stupid or didn't study. it's an open note quiz. i've been sick for a week and haven't learned any of this. @tkhunny

OpenStudy (anonymous):

What is the simplest form of the number? 27 ^2/3 (reads twenty seven to the two-thirds power) a) 3 b) 6 c) 9 d) 18 I think it's 18? Is that right? It's either 18 or 9 I'm pretty sure.

OpenStudy (anonymous):

 anyone know how to solve this one? lol.

OpenStudy (tkhunny):

Just answer the questions. \(\sqrt{x^{2}} = |x|\) \(\sqrt{x^{3}} = x\sqrt{x}\)